# A Continuous Energy Ising Machine Leveraging Difference-of-Convex Programming

### v1

@@ 5 Supplementary Materials
$\sum_{k=k_{0}}^{k_{1}-1}\delta_{k+1}\leqslant\frac{2\sigma}{\mu}\varphi(\Delta_{k_{0}})+\delta_{k_{0}}<\frac{r}{2}.$
Thus,
@@
which contradicts our assumption about $\boldsymbol{x}_{k_{1}}$. Thus, we have shown that $\|\boldsymbol{x}^{(k)}-\boldsymbol{x}^{*}\|_{2}<r$ for all $k>k_{0}$.
Similar to (S36), we have for all $K>k_{0}$ that
@@
This implies that the sequence $\{\boldsymbol{x}^{(k)}\}$ is Cauchy in $\mathbb{R}^{n}$ and is therefore convergent. However, $\boldsymbol{x}^{*}$ being a limit point of $\{\boldsymbol{x}^{(k)}\}$, the entire sequence must converge to $\boldsymbol{x}^{*}$. ∎
Unlike DOCH, providing convergence guarantees for its accelerated variant, ADOCH, is more challenging. This difficulty arises primarily from the use of the Barzilai-Borwein-type update [55] in ADOCH, which can lead to nonmonotonic behaviour in the objective sequence $\{\mathcal{H}(\boldsymbol{x}^{(k)})\}$. On the other hand, this scheme allows for more effective exploration of the energy landscape, often leading to better local minima of $\mathcal{H}$. Empirically, we found that ADOCH gives higher-quality ground states than DOCH, particularly for large Ising models. We wish to clarify that even if we assume that the ADOCH iterates $\{\boldsymbol{x}^{(k)}\},\{\boldsymbol{y}^{(k)}\}$ are bounded, it is difficult to guarantee convergence of objective values $\{\mathcal{H}(\boldsymbol{x}^{k})\}$ to a local minimum of $\mathcal{H}$ [41].
@@
The DOCH iterates $\{\boldsymbol{x}^{(k)}\}$ are bounded for any $\alpha>0$ provided
@@
The matrix norm $\|\cdot\|_{\infty}$ is induced by the vector norm $\|\boldsymbol{z}\|_{\infty}=\max_{1\leqslant i\leqslant n}\,|z_{i}|$. Therefore, by the definition of induced norm,
@@
We have seen that the DOCH iterates converge to a critical point of $\mathcal{H}$. Under additional assumptions, the critical point is guaranteed to be a local minimizer.
@@
To prove that $\boldsymbol{x}^{*}$ is a strict minimizer of $\mathcal{H}$, it suffices to show that the Hessian $\nabla^{2}\mathcal{H}(\boldsymbol{x}^{*})$ is positive definite. Now, it follows from definition (S9) that
@@
Comparing (S41) and (S43), we see that the condition $\|\mathbf{J}_{\mathcal{T}}(\boldsymbol{x}^{*})\|_{2}<1$ implies $\nabla^{2}\mathcal{H}(\boldsymbol{x}^{*})$ being positive definite. This establishes our claim. ∎
@@
In practice, the fixed points of $\mathcal{T}$ have nonzero components, which are rounded to get the spin vector. If these fixed points are stable in the sense of (S40), then Proposition 11 ensures that the iterates $\{\boldsymbol{x}^{(k)}\}$ converge to a local minimum of the Hamiltonian (see Figure S1). Empirically, we found that setting $\beta$ sufficiently large—on the order of $\mathcal{O}\left(n\sqrt{n}\|\mathbf{J}+\alpha\mathbf{I}\|_{\infty}\right)$, as stipulated by Proposition 10—keeps the Jacobian norm at the fixed point below one (see Figure S2). We found that this also yields better quality solutions (see Figure S3). In summary, this means that we can set $\beta$ to a high value and tune $\alpha$ freely.
Figure S1: Log-log plot of convergence behaviour for the DOCH algorithm. The plot shows the norm of the difference of successive iterations with iterations. A $100$-spin SK model is considered. We observe rapid convergence of the iterates $\{\boldsymbol{x}^{(k)}\}$ toward a fixed point $\boldsymbol{x}^{*}$ within 100 iterations, as computed via the fixed-point update rule (S38). Results are shown for various combinations of hyperparameters: $\alpha_{r}\in\{0.50,1.50\}$ and $\beta_{r}\in\{1,5,10\}$.
Figure S2: Dependence of the (average) norm of the Jacobian $\|\mathbf{J}_{\mathcal{T}}(\boldsymbol{x}^{*})\|_{2}$ on the ratio $\alpha_{r}$. We consider a $100$-spin SK model. The ratios are defined as $\alpha_{r}=\alpha/\lambda_{\max}(-\mathbf{J})$ and $\beta_{r}=\beta/(n\sqrt{n}\|\mathbf{J}+\alpha\mathbf{I}\|\infty)$. For each $\alpha_{r}\in(0,2)$ and fixed $\beta_{r}\in\{1,5,10\}$, we compute the fixed point $\boldsymbol{x}^{*}$ by applying the update rule in (38) for 100 iterations. The spectral norm of the Jacobian $\|\mathbf{J}_{\mathcal{T}}(\boldsymbol{x}^{*})\|_{2}$ is then evaluated at the resulting point. This process is repeated over multiple random initializations $\boldsymbol{x}^{(0)}$ and the Jacobian norms are averaged.
@@
We show that for certain parameter settings and under some assumptions, the Hamiltonians in [44, 26] resemble our Hamiltonian. The Hamiltonian for the Kerr-nonlinear parametric oscillator (KPO) is of the form
$\mathcal{H}_{c}(\boldsymbol{x},\boldsymbol{y},t)=\sum_{i=1}^{N}\left(\frac{K}{4}(x_{i}^{2}+y_{i}^{2})^{2}-\frac{p(t)}{2}(x_{i}^{2}-y_{i}^{2})+\frac{\Delta_{i}}{2}(x_{i}^{2}+y_{i}^{2})\right)-\frac{\xi_{0}}{2}\sum_{i=1}^{N}\sum_{j=1}^{N}J_{i,j}(x_{i}x_{j}+y_{i}y_{j})$ (S44)
@@
$\mathcal{H}_{KPO}(\boldsymbol{x},\boldsymbol{y},t)$ $=\sum_{i=1}^{N}\left(\frac{K}{4}x_{i}^{4}-\frac{p(t)-\Delta_{i}}{2}x_{i}^{2}\right)-\frac{\xi_{0}}{2}\sum_{i=1}^{N}\sum_{j=1}^{N}J_{i,j}x_{i}x_{j}$
@@
We explain how the MAX-CUT problem on a graph can be formulated as an Ising model and how the Goemans-Williamson Semidefinite Program (GW-SDP) [39] can be used to approximate its ground state.
@@
We can solve (S48) in polynomial time using standard interior-point methods. However, the computational complexity for achieving a given accuracy is relatively high, namely $\mathcal{O}(n^{4.5})$ for an n-spin Ising model [43]. In practice, this limits scalability to about $10^{3}$ spins. It was famously shown in [39] that by using a randomized rounding scheme to extract a binary spin $\boldsymbol{s}\in\{-1,1\}^{n}$ from the optimal solution of (S48), the expected cut value is at least $87.8\%$ of the optimal value of (S47).
@@
For small matrices, the largest eigenvalue $\lambda_{\max}(\mathbf{J})$ is computed using the power method [61]. For $n\geqslant 10^{4}$, we approximate $\lambda_{\max}(\mathbf{J})$ using Wigner’s semicircle law [48]: $\lambda_{\max}(\mathbf{J})\approx 2\langle\mathbf{J}\rangle\sqrt{n}$, where $\langle\mathbf{J}\rangle$ is the empirical variance of the entries of $\mathbf{J}$ (see main text).
@@
The look-back parameter $q$ in ADOCH is assigned a value between 5 and 10, depending on the structure of the Ising model. For larger Ising models $n\geqslant 10^{4}$, we take $q=5$.
@@
We implemented the Simulated Annealing (SA) algorithm following the approach in [31]. The pseudocode is provided below. We used a logarithmic cooling schedule $\beta(t)=\beta_{0}\log\,(1+t/T)$, where $\beta_{0}$ is the initial temperature and $T$ is the total number of iterations. SA has two tunable parameters, $\beta_{0}$ and $T$. For the $K_{2000}$ Ising model, the ratio $T/\beta_{0}\approx 1000$ is recommended in [31]; accordingly, we use $\beta_{0}=1$ and $T=1000$. For other cases, $\beta_{0}$ is selected from the interval $[1,2]$ to obtain the best performance.
Algorithm S1 Simulated Annealing
@@
9:     randomly sample $z$ from (0, 1)
@@
We implemented the ballistic Bifurcation Machine following the approach described in [27]. The pseudocode is provided below. Following the setup in [27], the pump rate is set to $a_{0}=1$, so that $a_{t}$ increases linearly from 0 to 1 according to the prescribed pump rate schedule. The time step $\Delta_{t}$ is fixed at 1 for the $K_{2000}$ and SK models. For MAX-CUT and large graphs, we select $\Delta_{t}$ from the set {0.25, 0.5, 0.75, 1, 1.25}, based on empirical performance. The coupling strength $c_{0}$ is set using the formula:
@@
Algorithm S2 ballistic Simulated Bifurcation Machine
@@
5: for $t=1$ to $T$ do
6:     $a_{t}\leftarrow\frac{a_{0}t}{T}$ $\triangleright$ pump rate schedule
@@
We have implemented the Simulated Coherent Ising Machine following [31]. The pseudocode is provided below, where $\mathcal{N}[\mathbf{0},\mathbf{I}]$ is the standard $n$-dimensional Gaussian distribution. Since the structure of the update rule closely resembles that of the Simulated Bifurcation algorithm [27], we adopt the same parameter settings for the pump rate $a_{0}$, the coupling strength $c_{0}$, and the time step $\Delta_{t}$. The noise amplitude A is chosen from the interval $[0.1,1]$, with values in this range empirically yielding the best performance.
Algorithm S3 Simulated Coherent Ising Machine
@@
7:     $\boldsymbol{x}\leftarrow\boldsymbol{x}+\left(-(a_{0}-a_{t})\boldsymbol{x}+c_{0}\mathbf{J}\,\mathrm{sign}(\boldsymbol{x})\right)dt+A\boldsymbol{w}\sqrt{\Delta_{t}}$
@@
We implemented the Spring Ising Algorithm (SIA) as described in [38]. The corresponding pseudocode is provided below. The function $\mathrm{Boundary}(\boldsymbol{q},\boldsymbol{p})$ enforces the following componentwise constraints on the vectors $\boldsymbol{q}$ and $\boldsymbol{p}$:
Following [38], we set the mass coefficient $m=1$, the elastic coefficient $k=0.5$, and the scaling coefficient $\zeta(t)$ to vary linearly from $0.8\zeta_{0}$ to $10\zeta_{0}$, where $\zeta_{0}=0.05$ is the base value. The time step $\Delta$ is selected within the interval $(0,1]$ for the best result. The momentum vector $\boldsymbol{p}$ is initialized such that each component $p_{i}$ is sampled uniformly from the range (0.0005,0.0005)(-0.0005,0.0005).
Algorithm S4 Spring Ising Algorithm
@@
Supplementary Table S1: Ising model parameters used for benchmarking.
@@
To benchmark our algorithm, we construct large to ultra-large coupling matrices $\mathbf{J}$, in both dense and sparse forms. A sparse $\mathbf{J}$ has approximately $(1-p)\%$ nonzero entries. We generate only the lower triangular part of the matrix $\mathbf{J}$, excluding the diagonal (which has zeros), and then symmetrize it by mirroring the values to the upper triangle. For the lower triangular part, we sample $z\in\{1,\dots,N_{p}\}$, with $N_{p}=\lfloor 102300/p\rfloor$; if $z<1023$, we set $J_{ij}=J_{ji}=z-511$. To manage large matrices ($n\geqslant 10^{5}$), we use the Compressed Sparse Row (CSR) format [62], which stores only nonzero values, column indices, and row offsets. This enables efficient storage and access. Using CSR, we generate sparse matrices with $n=10^{5}$ to $10^{8}$, and connectivity levels from $10^{-6}$ to $10^{-1}$, yielding billions of nonzeros with a fraction of the memory required by dense formats.
For ultra-large fully-connected coupling matrices, explicit storage is prohibitively expensive due to the quadratic memory requirement. To address this, we implement an efficient, memory-free approach based on procedural generation. Instead of considering the full $n\times n$ matrix, we use a deterministic pseudo-random function to generate matrix elements on demand. Specifically, each coupling coefficient is generated using the rule: $J_{ij}=\sin(i\,j+\mathrm{seed})$ with $\mathrm{seed}=100$. This function-based representation ensures that the matrix remains symmetric and reproducible while entirely eliminating the need to store it in memory. As a result, we are able to work with fully-connected matrices of size exceeding $10^{5}\times 10^{5}$, which would otherwise require terabytes of memory if stored explicitly.

### v2

@@ 5 Supplementary Material
$\sum_{k=k_{0}}^{k_{1}-1}\delta_{k+1}\leqslant\frac{2\sigma}{\mu}\varphi(\Delta_{k_{0}})+\delta_{k_{0}}<\frac{r}{2},$
and
@@
which contradicts our assumption about $\boldsymbol{x}_{k_{1}}$. Thus, must we have $\|\boldsymbol{x}^{(k)}-\boldsymbol{x}^{*}\|_{2}<r$ for all $k>k_{0}$.
Similar to (S37), for all $K>k_{0}$, we have
@@
Thus the sequence $\{\boldsymbol{x}^{(k)}\}$ is Cauchy and therefore convergent. Since $\boldsymbol{x}^{*}$ is a limit point of $\{\boldsymbol{x}^{(k)}\}$, it follows that the entire sequence converges to $\boldsymbol{x}^{*}$. ∎
Unlike DOCH, providing a convergence guarantee for ADOCH is more challenging. This difficulty arises primarily from the use of the Barzilai-Borwein-type update [grippo2002] in ADOCH, which can lead to nonmonotonic behaviour in the objective sequence $\{\mathcal{H}(\boldsymbol{x}^{(k)})\}$. On the other hand, this scheme allows for more effective exploration of the energy landscape, often leading to better local minima of $\mathcal{H}$. Empirically, we found that ADOCH gives higher-quality ground states than DOCH, particularly for large Ising models. We wish to clarify that even if we assume that the ADOCH iterates $\{\boldsymbol{x}^{(k)}\},\{\boldsymbol{y}^{(k)}\}$ are bounded, it is difficult to guarantee convergence of objective values $\{\mathcal{H}(\boldsymbol{x}^{k})\}$ to a local minimum of $\mathcal{H}$ [phan2018].
@@
The DOCH iterates $\{\boldsymbol{x}^{(k)}\}$ are bounded if $\alpha>0$ and
@@
The matrix norm $\|\cdot\|_{\infty}$ is induced by the vector norm $\|\boldsymbol{z}\|_{\infty}=\max_{1\leqslant i\leqslant n}\,|z_{i}|$. Hence,
@@
We have established that the DOCH iterates converge to a critical point of $\mathcal{H}$. Under additional assumptions, this critical point is in fact guaranteed to be a local minimizer.
@@
To prove that $\boldsymbol{x}^{*}$ is a strict minimizer of $\mathcal{H}$, it suffices to show that the Hessian $\nabla^{2}\mathcal{H}(\boldsymbol{x}^{*})$ is positive definite. It follows from definition (S11) that
@@
Comparing (S42) and (S44), we see that the condition $\|\mathbf{J}_{\mathcal{T}}(\boldsymbol{x}^{*})\|_{2}<1$ implies that $\nabla^{2}\mathcal{H}(\boldsymbol{x}^{*})$ is positive definite. This establishes our claim. ∎
@@
In practice, the fixed points of $\mathcal{T}$ have nonzero components, which are rounded to get the spin vector. If these fixed points are stable in the sense of (S41), then Proposition 11 ensures that the iterates $\{\boldsymbol{x}^{(k)}\}$ converge to a local minimum of the Hamiltonian (see Supplementary Figure S15). Empirically, we found that setting $\beta$ sufficiently large, on the order of $\mathcal{O}\left(n\sqrt{n}\|\mathbf{J}+\alpha\mathbf{I}\|_{\infty}\right)$, as stipulated by Proposition 10, keeps the Jacobian norm at the fixed point below one (see Supplementary Figure S16). We found that this also yields better quality solutions (see Supplementary Figure S17). In summary, this means that we can set $\beta$ to a high value and tune $\alpha$ freely.
Figure S15: Log-log plot of convergence behaviour for the DOCH algorithm. The plot shows the norm of the difference of successive iterates with iteration count. A $100$-spin SK model is considered. We observe rapid convergence of the iterates $\{\boldsymbol{x}^{(k)}\}$ toward a fixed point $\boldsymbol{x}^{*}$ within $100$ iterations, as computed via the fixed-point update rule (S39). Results are shown for various combinations of hyperparameters: $\alpha_{r}\in\{0.50,1.50\}$ and $\beta_{r}\in\{1,5,10\}$.
Figure S16: Dependence of the (average) norm of the Jacobian $\|\mathbf{J}_{\mathcal{T}}(\boldsymbol{x}^{*})\|_{2}$ on the ratio $\alpha_{r}$. We consider a $100$-spin SK model. The ratios are defined as $\alpha_{r}=\alpha/\lambda_{\max}(-\mathbf{J})$ and $\beta_{r}=\beta/(n\sqrt{n}\|\mathbf{J}+\alpha\mathbf{I}\|\infty)$. For each $\alpha_{r}\in(0,2)$ and fixed $\beta_{r}\in\{1,5,10\}$, we compute the fixed point $\boldsymbol{x}^{*}$ by applying the update rule in (S39) for $100$ iterations. The spectral norm of the Jacobian $\|\mathbf{J}_{\mathcal{T}}(\boldsymbol{x}^{*})\|_{2}$ is then evaluated at the resulting point. This process is repeated over multiple random initializations $\boldsymbol{x}^{(0)}$ and the average Jacobian norms are plotted against $\alpha_{r}$.
@@
Remarks 2.
We can leverage this convergence guarantee to introduce a principled stopping rule. One option is to monitor the relative error
$r^{(k)}=\frac{\|\boldsymbol{x}^{(k+1)}-\boldsymbol{x}^{(k)}\|_{2}}{\|\boldsymbol{x}^{(k)}\|_{2}},$
and stop once $r^{(k)}<\epsilon$, for some tolerance $\epsilon>0$. We tested this criterion on the K2000 graph (see Supplementary Figure S18), and found that it accurately detects the saturation point without any loss in performance. The other solvers lack this flexibility, as they lack comparable convergence guarantees and depend critically on cooling schedules.
Figure S18: Comparison of two different termination criteria: total runtime budget ($T=1$ s) vs. relative error tolerance ($\epsilon=10^{-3}$), for DOCH and ADOCH solvers on the K2000 graph. The results were obtained on a laptop equipped with a 4GB NVIDIA RTX 3050 GPU.
@@
We show that for certain parameter settings and under some assumptions, the Hamiltonians in [goto2018, goto2019] resemble our Hamiltonian. The Hamiltonian for the Kerr-nonlinear parametric oscillator (KPO) is of the form
$\mathcal{H}_{c}(\boldsymbol{x},\boldsymbol{y},t)=\sum_{i=1}^{N}\left(\frac{K}{4}(x_{i}^{2}+y_{i}^{2})^{2}-\frac{p(t)}{2}(x_{i}^{2}-y_{i}^{2})+\frac{\Delta_{i}}{2}(x_{i}^{2}+y_{i}^{2})\right)-\frac{\xi_{0}}{2}\sum_{i=1}^{N}\sum_{j=1}^{N}J_{ij}(x_{i}x_{j}+y_{i}y_{j})$ (S45)
@@
$\mathcal{H}_{KPO}(\boldsymbol{x},\boldsymbol{y},t)$ $=\sum_{i=1}^{N}\left(\frac{K}{4}x_{i}^{4}-\frac{p(t)-\Delta_{i}}{2}x_{i}^{2}\right)-\frac{\xi_{0}}{2}\sum_{i=1}^{N}\sum_{j=1}^{N}J_{ij}x_{i}x_{j}$
@@
We explain how the MAX-CUT problem on a graph can be reformulated as an Ising model and how the Goemans-Williamson Semidefinite Program (GW-SDP) [goemans1995] can be used to approximate its ground state.
@@
We can solve (S49) in polynomial time using standard interior-point methods. However, the computational complexity for achieving a given accuracy is relatively high, namely $\mathcal{O}(n^{4.5})$ for an n-spin Ising model [Luo2010]. In practice, this limits scalability to about $10^{3}$ spins. It was famously shown in [goemans1995] that by using a randomized rounding scheme to extract a binary spin $\boldsymbol{s}\in\{-1,1\}^{n}$ from the optimal solution of (S49), the expected cut value is at least $87.8\%$ of the optimal value of (S48).
@@
For small matrices, the largest eigenvalue $\lambda_{\max}(\mathbf{J})$ is computed using the power method [Horn_Johnson_1985]. For $n\geqslant 10^{4}$, we approximate $\lambda_{\max}(\mathbf{J})$ using Wigner’s semicircle law [mehta2004rmt]: $\lambda_{\max}(\mathbf{J})\approx 2\langle\mathbf{J}\rangle\sqrt{n}$, where $\langle\mathbf{J}\rangle$ is the empirical variance of the entries of $\mathbf{J}$ (see main text).
@@
The look-back parameter $q$ in ADOCH is assigned a value between 5 and 10, depending on the structure of the Ising model. For larger Ising models $n\geqslant 10^{4}$, we take $q\leqslant 5$.
@@
We implemented the Simulated Annealing (SA) algorithm following the approach in [inagaki2016]. The pseudocode is provided in Algorithm S1. We used a logarithmic cooling schedule $\beta(t)=\beta_{0}\log\,(1+t/T)$, where $\beta_{0}$ is the initial inverse temperature parameter and $T$ is the total number of iterations (or runtime). SA has two tunable parameters, $\beta_{0}$ and $T$. For the $K_{2000}$ Ising model, the ratio $T/\beta_{0}\approx 1000$ is recommended in [inagaki2016]; accordingly, we use $\beta_{0}=1$ and $T=1000$. For other cases, $\beta_{0}$ is selected from the interval $[1,2]$ to obtain the best performance.
Algorithm S1: Simulated Annealing
@@
9:     uniformly sample $z$ from (0, 1)
@@
We implemented the ballistic Bifurcation Machine following the approach described in [goto2021]. The pseudocode is provided in Algorithm S2. Following the setup in [goto2021], the pump rate is set to $a_{0}=1$, so that $a_{t}$ increases linearly from 0 to 1 according to the prescribed pump rate schedule. The time step $\Delta_{t}$ is selected from the set {0.25, 0.5, 0.75, 1, 1.25}, based on empirical performance. The coupling strength $c_{0}$ is set using the formula:
@@
Algorithm S2: ballistic Simulated Bifurcation Machine
@@
5: for $t=0$ to $T$ do
6:     $a_{t}\leftarrow(a_{0}t/T)$ $\triangleright$ pump rate schedule
@@
We implemented the Simulated Coherent Ising Machine following [goto2021]. The pseudocode is provided below, where $\mathcal{N}[\mathbf{0},\mathbf{I}]$ denotes the standard $n$-dimensional Gaussian distribution. We adopt the same parameter settings from [goto2021] for the pump rate $a_{0}$, the coupling strength $c_{0}$, and the time step $\Delta_{t}$. The noise amplitude $A$ is chosen from the interval $[0.1,1]$, with values in this range empirically yielding the best performance.
Algorithm S3: Simulated Coherent Ising Machine
@@
7:     $\boldsymbol{x}\leftarrow\boldsymbol{x}+\left(-(a_{0}-a_{t})\boldsymbol{x}+c_{0}\mathbf{J}\,\mathrm{sign}(\boldsymbol{x})\right)\Delta_{t}+A\boldsymbol{w}\sqrt{\Delta_{t}}$
@@
The original Coherent Ising Machine (CIM) [inagaki2016] is based on a degenerate optical parametric oscillator (DOPO) and is modeled by the following system of stochastic differential equations [inagaki2016, wang2013coherent]:
$dc_{i}=\bigg[(-1+p-c_{i}^{2}-s_{i}^{2})c_{i}+r\sum_{j}J_{ij}\tilde{c}_{j}\bigg]dt+\frac{1}{A_{s}}\sqrt{c_{i}^{2}+s_{i}^{2}+0.5}\,dW_{1},$ (S51)
$ds_{i}=\left[(-1-p-c_{i}^{2}-s_{i}^{2})s_{i}\right]dt+\frac{1}{A_{s}}\sqrt{c_{i}^{2}+s_{i}^{2}+0.5}\,dW_{2},$ (S52)
where
  • $c_{i}$ and $s_{i}$ are the in-phase and quadrature amplitudes of the $i$-th DOPO,
  • $\tilde{c}_{i}$ is the measured in-phase amplitude (perturbed by measurement noise),
  • $\{J_{ij}\}$ are the the coupling coefficients with coupling strength $r$,
  • $p$ is the normalized pump rate,
  • $A_{s}$ is the saturation amplitude,
  • $dW_{1}$ and $dW_{2}$ are independent Gaussian Wiener processes representing vacuum and pump fluctuations for in-phase and quadrature components, respectively.
In Supplementary Figure S19, we simulate the CIM dynamics for the MAX-CUT problem on the K2000 graph. To numerically solve (S51) and (S52), we apply the Euler-Maruyama integrator [maruyama1955euler], yielding the updates
$c_{i}(t+1)=c_{i}(t)+\bigg[(-1+p(t)-c_{i}(t)^{2}-s_{i}(t)^{2})c_{i}(t)\,+\,$ $r\sum_{j}J_{ij}\widetilde{c}_{j}(t)\bigg]\Delta_{t}$
$+\frac{w_{i}\sqrt{\Delta_{t}}}{A_{s}}\sqrt{c_{i}(t)^{2}+s_{i}(t)^{2}+0.5},$ (S53)
and
$s_{i}(t+1)$ $=s_{i}(t)+\Big[\left(-1-p(t)-c_{i}(t)^{2}-s_{i}(t)^{2}\right)s_{i}(t)\Big]\Delta_{t}+\frac{w_{i}^{\prime}\sqrt{\Delta_{t}}}{A_{s}}\sqrt{c_{i}(t)^{2}+s_{i}(t)^{2}+0.5}.$ (S54)
We have used $\widetilde{c}_{i}=c_{i}$ and $\widetilde{c}_{i}=\mathrm{sign}(c_{i})$ (clamped) in the experiments. The pump rate is defined as $p(t)=t/T$, and $w_{i},w_{i}^{\prime}$ are sampled independently from the standard normal distributions. After runtime $T$, the $i$-th spin state is obtained by taking the $\mathrm{sign}$ of $c_{i}(T)$. The hyperparameters $r,A_{s},\Delta_{t}$ are tuned for the best results. In Supplementary Figure S19, we observe that Algorithm S3 for SimCIM (adapted from [goto2021]) outperforms the simulated version of CIM in [inagaki2016] based on (S51) and (S52).
Figure S19: Comparison of simulated CIM [inagaki2016] with SimCIM. Histogram plots of cut values obtained for K2000 graph, with $100$ different initializations and two runtimes ($T=5$ ms, $50$ ms). (a), (b) Cut values obtained by two variants of simulated CIM (from equations (5), (S54)). (c) Cut values obtained by SimCIM (from Algorithm S2). The results were obtained on a laptop equipped with a 4GB NVIDIA RTX 3050 GPU.
@@
We implemented the Spring Ising Algorithm (SIA) as described in [jiang2024]. The corresponding pseudocode is provided in Algorithm S4. Following [jiang2024], we set the mass coefficient $m=1$, the elastic coefficient $k=0.5$, and the scaling coefficient $\zeta(t)$ to vary linearly from $0.8\zeta_{0}$ to $10\zeta_{0}$, where $\zeta_{0}=0.05$ is the base value. The time step $\Delta$ is selected within the interval $(0,1]$ for the best result. The momentum vector $\boldsymbol{p}$ is initialized such that each component $p_{i}$ is sampled uniformly from the range (0.0005,0.0005)(-0.0005,0.0005).
Algorithm S4: Spring Ising Algorithm
@@
The function $\mathrm{Boundary}(\boldsymbol{q},\boldsymbol{p})$ enforces the following componentwise constraints on the vectors $\boldsymbol{q}$ and $\boldsymbol{p}$:
@@
Table S1: Ising model parameters used for benchmarking.
@@
To benchmark our algorithm, we construct large to ultra-large coupling matrices $\mathbf{J}$, in both dense and sparse forms. A sparse $\mathbf{J}$ has approximately $(1-p)\%$ nonzero entries. We generate only the lower triangular part of the matrix $\mathbf{J}$, excluding the diagonal (which has zeros), and then symmetrize it by mirroring the values to the upper triangle. To sample $p$-sparse matrix $\mathbf{J}$ with $J_{ij}$ from $9$-bit signed integers, we sample $z\in\{1,\dots,N_{p}\}$, with $N_{p}=\lfloor 102300/p\rfloor$; if $z<1023$, we set $J_{ij}=J_{ji}=z-511$ (described in detail in Algorithm S5). To manage large matrices ($n\geqslant 10^{5}$), we use the Compressed Sparse Row (CSR) format [CSR2003book], which stores only nonzero values, column indices, and row offsets. This enables efficient storage and access. Using CSR, we generate sparse matrices with $n=10^{5}$ to $10^{8}$, and connectivity levels from $10^{-6}$ to $10^{-1}$, yielding billions of nonzeros with a fraction of the memory required by dense formats.
For ultra-large dense coupling matrices, explicit storage is prohibitively expensive due to the quadratic memory requirement. To address this, we implement an efficient, memory-free approach based on procedural generation. Instead of considering the full $n\times n$ matrix, we use a deterministic pseudo-random function to generate matrix elements on demand. Specifically, each coupling coefficient is generated using the rule: $J_{ij}=\sin(i\,j+\mathrm{seed})$ with $\mathrm{seed}=100$. This function-based representation ensures that the matrix remains symmetric and reproducible while eliminating the need to store it in memory. As a result, we can work with dense matrices of size exceeding $10^{5}\times 10^{5}$, which would otherwise require terabytes of memory if stored explicitly.